There are 9 players on a softball team. How many ways can the coach pick the first 5 batters?
A 5
B 45
C 15,120
D 72,876

Answer by stanbon(75887)

There are 9 players on a softball team. How many ways can the coach pick the first 5 batters?
------------
9P5 = 9! / [9-5]! = 9*8*7*6*5=15120 ways
Cheers,
Stan H.